OpenStudy (anonymous):

Solve 3|2x + 2| − 7 = 5 for x. Show your work.

OpenStudy (anonymous):

start with \[3|2x+2|=12\] then \[|2x+2|=4\] then solve two equations \[2x2=4\] and \[2x+2=-4\] separately, get two different solutions

myininaya (myininaya):

so we have \[6|x+1|-7=5\] Add 7 on both sides \[6|x+1|=12 =>|x+1|=\frac{12}{6}=>|x+1|=2\]

myininaya (myininaya):

Now we have to solve \[x+1=2 \text{ or } x+1=-2\]

myininaya (myininaya):

I think satellite made a type-0 is first equation should be 2x+2=4 lol

myininaya (myininaya):

i made a type-o his* not is

OpenStudy (anonymous):

Distribute 3 to 2x+2. That equals 6x +6. 6x+6-7=5 Isolate the variable by moving all of the numbers over to one side. -6 -6 +7 +7 6x=6 Divide by 6 to get x alone x=6

myininaya (myininaya):

you mean x=1 which is one solution

myininaya (myininaya):

there is another solution x=-3

OpenStudy (anonymous):

yeah that's what i meant! :) I'm not that great at math, so some things slip. Anyways, that's how I think you should do it, but I can't be a hundred percent sure.